OpenStudy (anonymous):

An electric scooter has a battery capable of supplying 120 Wh of energy. If friction forces and other losses account for 60.0% of the energy usage, what altitude change can a rider achieve when driving in hilly terrain, if the rider and scooter have a combined weight of 890 N? @Physics

OpenStudy (anonymous):

\[Energy\ in\ battery\ =120\ W.hr\]\[=120\times3600\ W.s\]\[Energy\ in\ battery\ =4.32\times 10^5\ J\] \[Work\ supplied\ by\ battery\ =\ 60\%\ of\ Energy\ in\ battery\]\[=0.6\times 4.32\times10^5\]\[=2.592\times10^5\ J\] \[Energy\ gained\ by\ object\ =\ Work\ supplied\ by\ battery\]\[Also,\ Energy\ gained\ by\ object\ =\ m.g.h\]\[\implies(890\ N)\times h=2.592\times10^5\ J\]\[h=2.91236\times 10^2\ m\]

OpenStudy (anonymous):

says the answer is wrong =( but it says its within 10% of the correct answer!